OpenStudy (grant330sims):

Simplify. ({1/36}- {1/x^2}) / ({1/6}+{1/x})

OpenStudy (anonymous):

(x+18)(x-12)/36x(x+6) This is the answer if I have the formula absolutely correct.

OpenStudy (anonymous):

If I have this wrong, I do not see where, but it is possible. I hope I helped :)

OpenStudy (grant330sims):

one minute. I dont think it is completely simplified.

OpenStudy (grant330sims):

and if it is, it's not one of the answers on this quiz...

OpenStudy (anonymous):

Hmm.... Hold on I will re check it.

OpenStudy (anonymous):

6x^2-1/x Here. You were right, It was an unfinished simplification. My bad :)

OpenStudy (grant330sims):

thats not it either..

OpenStudy (anonymous):

Well then the format must be wrong for this roblem becuase I'm pretty proficient with these things

OpenStudy (anonymous):

({1/36}- {1/x^2}) / ({1/6}+{1/x}) = ({1/6}-{1/x})

OpenStudy (anonymous):

\[a^2-b^2=(a+b)(a-b) \]

OpenStudy (anonymous):

\(a= \frac 16 \) \(b=\frac 1x \)

OpenStudy (anonymous):

FFM thank you :)

OpenStudy (grant330sims):

well, I'm not questioning your abilities to do this, but these were the answers I was given. a.6-x/6x b.6x/x-6 c.x-6/6x d.6x-6-x

OpenStudy (anonymous):

(x-6)/6x; c

OpenStudy (grant330sims):

sweet. Now how do you do that? Sorry I'm taking so long.

OpenStudy (anonymous):

I have given you sufficient hints. Please show some effort on your part :)